OpenStudy (matlee):

Can anyone help me with a big science project? ... 3+3

OpenStudy (anonymous):

3+3=6

OpenStudy (fairegaming):

3+3= Soccerball^2 Which adds to 69 Soccerballs / 2 which is (34.5)^2

OpenStudy (matlee):

Illumanati confirmed

OpenStudy (matlee):

how does 3+3 =6 i dont understand

OpenStudy (fairegaming):

Pls gimmi medal for smartness

OpenStudy (fairegaming):

and help me w/ my question

OpenStudy (matlee):

Wait you must answer 1 lat question 1+1

OpenStudy (anonymous):

3+3=5 idots smh -_-

OpenStudy (matlee):

I knew it was 5 thats what my teacher said

OpenStudy (fairegaming):

1 + 1 = 21$ which adds to ^%$#^#1

OpenStudy (matlee):

Omg! She said that too. THANKS! Can you please trade each other medals

OpenStudy (matlee):

Whats your question